# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 4 Methodology
This section describes the approach used to carry out a Systematic Literature Review (SLR) with the goal of investigating Zero Trust architecture (ZTA) in light of its applications, issues and challenges, and enabling technologies. An SLR involves a methodical and structured approach to identifying, assessing, and interpreting relevant existing research with an aim to answer some research questions [59]. The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) reporting criteria is followed in this study. PRISMA framework is widely used in the literature and it provides a structured approach for conducting an SLR in a reproducible and transparent manner. Consequently, the methodologies employed to carry out the SLR are highlighted in this section. It comprises the questions that the study seeks to answer, the database from which the data is obtained, the search strategy, the procedure for screening and selecting the articles based on inclusion and exclusion criteria, and the study selection and characteristics.
@@ 4.2 INFORMATION SOURCES/DATABASES
Although publications such as journal articles, conference papers, review/survey papers, etc can be found in a variety of databases, three popular databases — Scopus, Web of Science (WoS), and IEEE Xplore — were chosen as information sources for this study. Their extensive article collection, general recognition and reputation in the academic community, and efforts to reduce the quantity of duplicate articles are the reasons behind this. The databases were thoroughly searched to find relevant articles for the study.
@@ 4.3 SEARCH STRATEGY/SEARCH TERMS
This step represents a crucial phase in conducting this study as it describes the strategy employed to identify the research articles reviewed based on the questions and objectives of the research. The selected databases differ slightly in the way they are queried. In essence, while publications can be searched in the Scopus database through their ”Title-Abstract-Keyword” together, these are seperate in WoS database and they have to be combined if needed. Therefore, in WoS database we repeated the search terms to search for the articles through the title and abstract with assumption to get the same effect and results as Scopus database. The following are the respective search term/string used in each of the databases to retrieve the relevant articles:
@@
While, in Web of Science (WoS), we used:
@@
The search progresses through several stages. Initially, we query the databases using the search terms only without applying any constraints. This yields 1,122, 539 and 642 articles from Scopus, WoS, and IEEE Xplore databases, respectively. We limit the search results by applying some filters, such as focusing on publications that reached final, written in English, from Computer Science and Engineering research area, while including only conference, article, and review papers . As a result, the number of retrieved articles is reduced to 839 articles in Scopus database, 485 articles in WoS database, and 458 in IEEE Xplore database. Finally, we restricted the search within a span of 10 years period from 2016 to 2025. This slightly reduces the results to 836, 482, and 456 articles from the Scopus, WoS, and IEEE Xplore databases, respectively. We believe the search strategy including the search keywords employed should capture all relevant publications to help us successfully answer our research questions.
@@ 4.5 STUDY SELECTION
The four step process of the PRISMA framework which are Identification, Screening, Eligibility, and Inclusion were used to select the articles for this study. Initially, after identifying and retrieving the research articles which resulted in a total of 1,774 research papers, 831 duplicate entries were eliminated using a Python script, 40 invalid entries usually from Scopus database, and 5 not available articles, resulting in 943 remaining articles. Following this, the artilces were screened in the second stage based on their titles and abstracts. At this stage, 522 articles were removed by title, and 278 artilces by abstract. 98 articles made it to the eligibility stage where we used full-text reading to assess them based on the inclusion and exclusion criteria and the research questions. Seven unrelated research articles were removed, leaving a total of 91 articles which were included in the SLR in the last stage of the process. These articles represent the primary source of information for the SLR.
@@ 4.6 Study Characteristics
This section describes some common characteristics of the studies included in this SLR. As depicted in Figure  5, 50% of the articles were journal publications, reflecting a strong preference for journals as the primary medium for publishing research on ZTA. , conference papers accounted for 42%, while review/survey papers represented 8%, indicating limited research on comprehensive reviews or surveys in the ZTA domain.
@@
In addition, Table  3 presents the distribution of the articles based on source titles, however, only sources with more than one publication are shown. Among these, IEEE Access made the largest contribution with 12 publications, representing approximately 13% of the reviewed studies. The IEEE Global Communications Conference, GLOBECOM has five conference papers, followed by Electronics (Switzerland) and Ad Hoc Networks journals having four and three articles, respectively. Other sources, including IEEE Wireless Communications, Computer Networks, Applied Sciences (Switzerland), Proceedings of IEEE/ACS International Conference on Computer Systems and Applications, AICCSA, and the Wireless Communications and Mobile Computing, each contributed two articles, collectively accounting for 11% of the total. The remaining 62% of the articles were distributed across a diverse set of journals and conferences, reflecting the broad application and multidisciplinary nature of Zero Trust Architecture.
@@ 5 Taxonomy
In this section we describe the proposed ZTA taxonomy developed in this SLR. Taxonomy is a systematic organization of components, concepts, or entities in a field or domain based on their shared characteristics and relationships. Thus, the ZTA taxonomy depicted in Figure  6 presents a structured way to explore ZTA by facilitating the identification of its various application domains, requirements, and enabling technologies. This classification not only simplifies the implementation process but also supports researchers and practitioners in analyzing ZTA across varying contexts.
@@ 5.1 Application Domains
Based on the existing literatures studied, it is evident that Zero Trust Architecture is being implemented across various sectors, each with unique security needs. In healthcare, the ZTA enables organizations to safeguard their patients’ sensitive medical records and health data with robust access controls, secure data exchange protocols, and real-time monitoring of electronic health records (EHRs). ZTA in the finance sector helps to combat cyber threats like financial fraud through encrypted communications, strong identity verification, and network micro-segmentation. For the Internet of Things (IoT) including Industrial IoT, ZTA addresses the inherent vulnerabilities of interconnected devices by establishing secure identities, managing access permissions, and enforcing trust evaluations. In cloud computing, ZTA supports multi-tenant environments with dynamic access policies, real-time authentication, and precise micro-segmentation, safeguarding data and applications in shared infrastructures. Likewise, as the 5G and 6G networks continue to evolve, ZTA offers greater security with continuous monitoring, dynamic trust evaluation, and distributed security policies designed for scenarios with high-speed, low-latency connectivity. The following subsections give a detailed discussion on these and other applications which indicate the applicability and increasing prevalence of ZTA security strategies in enhancing security defenses across diverse application domains.
@@ 5.1.1 Healthcare and Medical
The healthcare sector is a critical domain that relies heavily on computing devices especially with the proliferation of Internet of Health Things (IoHT) and handles diverse types of people data [87]. This data includes the health records or status of patients, their personal information, as well as financial details such as credit card information. The extensive devices used in the healthcare system expand its attack surface, while hosting large volume of sensitive data make it a high target for cyber attacks that can be executed if the system is compromised which can include identity theft, damage to reputation, and financial fraud [88]. Example of cyberattacks in the healthcare domain is the Conti ransomware attack on Ireland’s Health Service Executive (HSE) which occurred in May 2021 [89]. The attack encrypted around 80% of the HSE’s data, disrupted operations nationwide, made essential services such as radiotherapy to be suspended, and the hospitals had to revert to paper-based systems. Moreover, attacks on identity theft and reputational damage in the healthcare systems include the NationsBenefits data breach in 2023 where sensitive personal data of over 3 million individuals including social security numbers were exposed [90], and the Medibank breach in Australia in October 2022 which resulted in a major public relations crisis [91]. These incidents highlights the critical need for robust cybersecurity measures like the ZTA in the healthcare sector to protect sensitive data and maintain trust.
@@
Healthcare systems are composed of diverse collection of devices including IoT devices, legacy systems, and unmanageable devices [84] whose interactions and inter-connectivity exceed beyond the local healthcare facility to the internet, for example, the cloud, exchanging private and control messages. While this presents so many advantages, it has equally expanded the attack surface [92]. This highlighted the need for a robust security to secure the healthcare systems. From the reviewed articles, Erik et al [23] conducted a research to assess the feasibility and practicality of adopting ZTA in healthcare institutions by comparing its perceived potentials against PBSM in terms of control, security, cost-effectiveness, risk, supportability, and operational aspects. They used a healthcare company based in the United States with a capacity of about 50K individuals as a case study. While qualitative analysis favor ZTA over PBSM due its increased security because of the architecture’s micro segmentation requirement, but there are some associated overheads in terms of cost and support requirements with ZTA implementation. The findings of the research emphasize a hybrid approach that incorporates elements of both ZTA and PBSM, adhering to the seven ZTA tenets outlined in NIST SP 800-207.
@@
Similarly, in [84], a framework to help small-to-medium sized healthcare organizations implement ZTA is presented which was developed and tested using Cisco Modelling Labs (CML). The framework comprises of five stages and utilizes a hybrid approach which enables the use of traditional network security practices like VPN while incorporating ZT concepts. It starts from basic security which implement basic security controls like changing default credentials, multi-factor authentication (MFA), and data encryption to final stage of defense in-depth which implement security measures like DNS sinkholing, secure internet access for BYOD and remote users, and behavior analysis for DMZ servers. While the main objective of the framework is to prevent lateral movement, the authors acknowledged that simulation performance results may differ in real-world implementations. According to [93], layering is the approach to be applied when it comes to securing the healthcare sector against security threats and vulnerabilities especially during crisis like Covid-19. They proposed the ”Control Systems Cyber Security Reference Architecture (RA) for Critical Infrastructure”, which can be used by healthcare organizations to improve their cybersecurity strategies based on the principles of Defense-in-Depth (DiD) and Zero Trust. The proposed architecture serves as a reference that can guide healthcare facilities in designing their cybersecurity policies and practices which will ensure the continuity of operations and the availability of essential services , which is crucial in healthcare settings.
@@
To secure 5G-enabled smart healthcare environments, a multi-layered ZTA framework is proposed in [94], and emphasizes continuous trust evaluation and proactive defense to ensure the confidentiality, integrity, and availability of sensitive medical data. It begins with user enrollment, where individuals provide basic identity details. However, trust is not granted at this initial stage. Instead, a one-time password (OTP) is sent to the user’s email address, requiring successful input for identity validation. Access to sensitive patient information requires an additional layer of verification, again utilizing an OTP as a safeguard. The multi-layered approach, with OTPs acting as security safeguards, aims to deter unauthorized access and ensure the integrity of patient data within the smart healthcare domain. Furthermore, a ZT framework for e-health environments is proposed in [92] based on user trust levels and sensitivity of resources. The authors defined a trust score that ranges from 0 to 100, where 0 is no trust and 100 represents complete trust, and sensitivity score that bears the same value range. Access to a resource is granted, or denied, or may require re-authentication based on the trust level of the user and the sensitivity of the resource requested. In this system, user trust may be reduced by the system if it identifies any access abnormalities through three analysis mechanisms - context-based trust which evaluates how the user interacts with the system, device trust which assesses the device used for access, and historical trust which analyzes past user behavior. The effectiveness of the model was demonstrated through four simulated scenarios.
@@
With the increasing drive towards digital transformation, patients health records have been progressively digitized which enables easy administration and exchange of patient’s records where it can be stored, retrieved, or modified seamlessly [95]. However, this has exposes the records to various security challenges which require adequate security solutions [96]. On this note, Sultana et al. [88] utilized the combined advantages of blockchain immutability and the added security of ZTA and proposed a framework for secure transfer and storage of medical data and image files in an electronic healthcare system. In their work, they highlighted that blockchain ensure the integrity of data while the ZTA provides encryption and ensure that only authenticated devices and users access the enterprise network. The proposed system is decentralized to eliminate single point bottleneck, used two-factor authentication to secure access to the proposed system’s web page, smart contract to assign roles to users based on their functions and privileges, while verifying the security status and parameters of any device requesting access to the system.
@@
Furthermore, a ZTA framework integrated into microservice architecture to address security challenges in EHR is proposed in [95]. It leverages several ZTA principles which highlights the importance of continuous verification, granular access controls, and a data-centric approach to protect sensitive health information. The integration of ZTA with microservice architecture allows each microservice to enforce its own security policies, creating isolated trust zones within the system, and thus, limits the impact of a security breach in one service on other parts of the system. The effectiveness of the framework is demonstrated through a case study which shows significant improvements in security posture while maintaining acceptable system performance and user experience.

### v2

@@ 4 Methodology
This section describes the approach used to carry out a Systematic Literature Review (SLR) with the goal of investigating Zero Trust Architecture (ZTA) in light of its applications, issues and challenges, and enabling technologies. An SLR involves a methodical and structured approach to identifying, assessing, and interpreting relevant existing research with an aim to answer some research questions [59]. The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) reporting criteria is followed in this study [60]. PRISMA framework is widely used in the literature and it provides a structured approach for conducting an SLR in a reproducible and transparent manner. Consequently, the methodologies employed to carry out the SLR are highlighted in this section. It comprises the questions that the study seeks to answer, the database from which the data is obtained, the search strategy, the procedure for screening and selecting the articles based on inclusion and exclusion criteria, and the study selection and characteristics.
@@ 4.2 Information Sources/Databases
Although publications such as journal articles, conference papers, review/survey papers, etc. can be found in a variety of databases, three popular databases — Scopus, Web of Science (WoS), and IEEE Xplore — were chosen as information sources for this study. Their extensive article collection, general recognition and reputation in the academic community, and efforts to reduce the quantity of duplicate articles are the reasons behind this. The databases were thoroughly searched to find relevant articles for the study.
@@ 4.3 Search Strategy/Search Terms
This step represents a crucial phase in conducting this study as it describes the strategy employed to identify the research articles reviewed based on the questions and objectives of the research. The selected databases differ slightly in the way they are queried. In essence, while publications can be searched in the Scopus database through their ”Title-Abstract-Keyword” together, these are separate in the WoS database and they have to be combined if needed. Therefore, in WoS database we repeated the search terms to search for the articles through the title and abstract with assumption to get the same effect and results as Scopus database. The following are the respective search terms/strings used in each of the databases to retrieve the relevant articles:
@@
While, in the Web of Science (WoS), we used:
@@
The search progresses through several stages. Initially, we query the databases using the search terms only without applying any constraints. This yields 1,122, 539 and 642 articles from Scopus, WoS, and IEEE Xplore databases, respectively. We limit the search results by applying some filters, such as focusing on publications that reached final, written in English, from the Computer Science and Engineering research area, while including only conference, article, and review papers . As a result, the number of retrieved articles is reduced to 839 articles in the Scopus database, 485 articles in WoS database, and 458 in IEEE Xplore database. Finally, we restricted the search within a span of 10 years period from 2016 to 2025. This slightly reduces the results to 836, 482, and 456 articles from the Scopus, WoS, and IEEE Xplore databases, respectively. We believe that the search strategy including the search keywords employed should capture all relevant publications to help us successfully answer our research questions.
@@ 4.5 Study Selection
The four step process of the PRISMA framework [60] which are Identification, Screening, Eligibility, and Inclusion were used to select articles for this study. Initially, after identifying and retrieving the research articles which resulted in a total of 1,774 research papers, 831 duplicate entries were eliminated using a Python script, 40 invalid entries usually from the Scopus database, and 5 not available articles, resulting in 943 remaining articles. Following this, the articles were screened in the second stage based on their titles and abstracts. At this stage, 522 articles were removed by title, and 278 articles by abstract. 98 articles made it to the eligibility stage where we used full-text reading to assess them based on the inclusion and exclusion criteria and the research questions. Seven unrelated research articles were removed, leaving a total of 91 articles which were included in the SLR in the last stage of the process. These articles represent the primary source of information for the SLR.
4.6 Quality Assessment
The quality of the included publications is examined, which involves evaluating their reliability, methodological rigor, and relevance to the SLR based on a formulated set of questions. This represents an important stage in the selection of the included publications and the questions were defined after being inspired by existing SLRs such as [61], [62], [63], and [64] because there are no general criteria to assess the quality of a study [63], [65]. For each of the questions, a quality score of 1 is assigned to a given study if the answer is ”Yes”, 0.5 for ”Partially”, and 0 for ”No”.
  • QA1: Does the study explicitly focus on ZTA or related concepts?
  • QA2: Is the study relevant to at least one of the research questions?
  • QA3: Does the study provide a clear research objective and research methodology?
  • QA4: Are the research methods well defined, reproducible, and rigorous?
  • QA5: Does the study include empirical analysis, case studies, or experimental validation?
  • QA6: Is there a clear statement of results or findings?
Based on these questions, a publication has a final quality score that ranges from 0 to 6, determined by adding its individual scores. This ensures the inclusion of relevant and high-quality publications in the review analysis.
@@ 4.7 Study Characteristics
This section describes some common characteristics of the studies included in this SLR. As depicted in Figure  5, journal publications accounted for 58% of the total publications, comprising 50% articles, which reflects a strong preference and high volume of research on ZTA in article form, while review papers represented 8%, indicating limited research on comprehensive reviews or surveys in the ZTA domain. On the other hand, conference papers accounted for 42% of the total publications considered in this study. The classification of the publications into articles, reviews, and conference papers follows the categorization used in the Scopus and Web of Science databases which is commonly used in the literature.
@@
In addition, Table  3 presents the distribution of the publications based on source titles, however, only sources with more than one publication are shown. Among these, IEEE Access made the largest contribution with 12 publications, representing approximately 13% of the reviewed studies. The IEEE Global Communications Conference, GLOBECOM has five conference papers, followed by Electronics (Switzerland) and Ad Hoc Networks journals having four and three articles, respectively. Other sources, including IEEE Wireless Communications, Computer Networks, Applied Sciences (Switzerland), Proceedings of IEEE/ACS International Conference on Computer Systems and Applications, AICCSA, and the Wireless Communications and Mobile Computing, each contributed two articles, collectively accounting for 11% of the total. The remaining 62% of the articles were distributed across a diverse set of journals and conferences, reflecting the broad application and multidisciplinary nature of Zero Trust Architecture.
4.8 Threat to Validity
Over the years, research on ZTA has progressively evolved, gathering contributions from researchers through various types of publications, including journal articles and conferences papers. Although efforts have been made to capture all relevant publications through a number of approaches, the limitations of the SLR have to be considered to recognize potential gaps and provide a basis for future research to address any overlooked aspects. Two key threats to the validity of this SLR can be identified: one is the potential search incompleteness, which can lead to bias in identifying and selecting the reviewed publications, and the other is related to the databases considered.
The search terms presented in Section 4.3 were used to identify and select the reviewed publications. It is possible that some studies that used alternative terminologies may not have been included. However, these terms were deemed representative, as they were derived from various research articles on ZTA and the search closely adhered to the PRISMA framework. In addition, well-defined inclusion and exclusion criteria that capture relevant publications while avoiding an overly broad scope were used. Similarly, the search is limited to only three databases (Scopus, Web of Science, and IEEE Xplore) which may have potentially excluded relevant publications from other sources. These databases were chosen because of their extensive collection, wide recognition and reputation in the academic community, and to limit the number of duplicated entries.
@@ 5 Taxonomy
In this section we describe the proposed ZTA taxonomy developed in this SLR. Taxonomy is a systematic organization of components, concepts, or entities in a field or domain based on their shared characteristics and relationships. Thus, the ZTA taxonomy depicted in Figure  6 has three main categories - Application Domains, Application Requirements, and Enabling Technologies. We derived these categories through a systematic thematic analysis (iterative coding process) of the existing literature on ZTA where key themes were extracted, categorized based on conceptual similarities or authors’ proposed applications, and then refined through expert validation.
@@
The taxonomy presents a structured way to explore ZTA by facilitating the identification of its various application domains, requirements, and enabling technologies. More specifically, given the diverse security challenges across industries, the application domains represent the various sectors where ZTA has been applied or proposed, capture real-world use cases, and contextualize how ZTA principles are adapted based on sector-specific constraints. On the other hand, the application requirements define the foundational elements necessary for implementing ZTA and identifying them will help to distinguish true ZTA implementations from partial adoptions or traditional PBSM. Finally, the enabling technologies encompass cutting-edge technologies and innovations that facilitate the implementation of ZTA and ensure its effectiveness in practice. We believe that this classification not only simplifies the implementation process, but also supports researchers and practitioners in analyzing ZTA across varying contexts.
@@ 5.1 Application Domains
Based on the existing literature studied, it is evident that Zero Trust Architecture is being implemented across various sectors, each with unique security needs. In healthcare, the ZTA enables organizations to safeguard their patients’ sensitive medical records and health data with robust access controls, secure data exchange protocols, and real-time monitoring of electronic health records (EHRs). ZTA in the finance sector helps to combat cyber threats like financial fraud through encrypted communications, strong identity verification, and network micro-segmentation. For the Internet of Things (IoT) including Industrial IoT, ZTA addresses the inherent vulnerabilities of interconnected devices by establishing secure identities, managing access permissions, and enforcing trust evaluations. In cloud computing, ZTA supports multi-tenant environments with dynamic access policies, real-time authentication, and precise micro-segmentation, safeguarding data and applications in shared infrastructures. Likewise, as the 5G and 6G networks continue to evolve, ZTA offers greater security with continuous monitoring, dynamic trust evaluation, and distributed security policies designed for scenarios with high-speed, low-latency connectivity. The following subsections give a detailed discussion on these and other applications which indicate the applicability and increasing prevalence of ZTA security strategies in enhancing security defenses across diverse application domains.
@@ 5.1.1 Healthcare and Medical
The healthcare sector is a critical domain that relies heavily on computing devices especially with the proliferation of Internet of Health Things (IoHT) and handles diverse types of people data [93]. This data includes the health records or status of patients, their personal information, as well as financial details such as credit card information. The extensive devices used in the healthcare system expand its attack surface, while hosting large volume of sensitive data make it a high target for cyber attacks that can be executed if the system is compromised which can include identity theft, damage to reputation, and financial fraud [94]. Example of cyberattacks in the healthcare domain is the Conti ransomware attack on Ireland’s Health Service Executive (HSE) which occurred in May 2021 [95]. The attack encrypted around 80% of the HSE’s data, disrupted operations nationwide, made essential services such as radiotherapy to be suspended, and the hospitals had to revert to paper-based systems. Moreover, attacks on identity theft and reputational damage in the healthcare systems include the NationsBenefits data breach in 2023 where sensitive personal data of over 3 million individuals including social security numbers were exposed [96], and the Medibank breach in Australia in October 2022 which resulted in a major public relations crisis [97]. These incidents highlight the critical need for robust cybersecurity measures like the ZTA in the healthcare sector to protect sensitive data and maintain trust.
@@
Healthcare systems are composed of diverse collection of devices including IoT devices, legacy systems, and unmanageable devices [90] whose interactions and interconnectivity extend beyond the local healthcare facility to the internet, for example, the cloud, exchanging private and control messages. While this presents so many advantages, it has equally expanded the attack surface [98]. This highlighted the need for a robust security to secure the healthcare systems. From the reviewed articles, Erik et al [23] conducted a research to assess the feasibility and practicality of adopting ZTA in healthcare institutions by comparing its perceived potentials against PBSM in terms of control, security, cost effectiveness, risk, supportability, and operational aspects. They used a healthcare company based in the United States with a capacity of about 50K individuals as a case study. While qualitative analysis favor ZTA over PBSM due its increased security because of the architecture’s micro-segmentation requirement, but there are some associated overheads in terms of cost and support requirements with ZTA implementation. The findings of the research emphasize a hybrid approach that incorporates elements of both ZTA and PBSM, adhering to the seven ZTA tenets outlined in NIST SP 800-207.
@@
Similarly, in [90], a framework to help small to medium-sized healthcare organizations implement ZTA is presented which was developed and tested using Cisco Modelling Labs (CML). The framework comprises of five stages and utilizes a hybrid approach which enables the use of traditional network security practices like VPN while incorporating ZT concepts. It starts from basic security which implement basic security controls like changing default credentials, multifactor authentication (MFA), and data encryption to final stage of defense in-depth which implement security measures like DNS sinkholing, secure internet access for BYOD and remote users, and behavior analysis for DMZ servers. While the main objective of the framework is to prevent lateral movement, the authors acknowledged that simulation performance results may differ in real-world implementations. According to [99], layering is the approach to be applied when it comes to securing the healthcare sector against security threats and vulnerabilities, especially during crises like Covid-19. They proposed the ”Control Systems Cyber Security Reference Architecture (RA) for Critical Infrastructure”, which can be used by healthcare organizations to improve their cybersecurity strategies based on the principles of Defense-in-Depth (DiD) and Zero Trust. The proposed architecture serves as a reference that can guide healthcare facilities in designing their cybersecurity policies and practices which will ensure the continuity of operations and the availability of essential services , which is crucial in healthcare settings.
@@
To secure 5G-enabled smart healthcare environments, a multilayered ZTA framework is proposed in [100], and emphasizes continuous trust evaluation and proactive defense to ensure confidentiality, integrity, and availability of sensitive medical data. It begins with user enrollment, where individuals provide basic identity details. However, trust is not granted at this initial stage. Instead, a one-time password (OTP) is sent to the user’s email address, requiring successful input for identity validation. Access to sensitive patient information requires an additional layer of verification, again utilizing an OTP as a safeguard. The multilayered approach, with OTPs acting as security safeguards, aims to deter unauthorized access and ensure the integrity of patient data within the smart healthcare domain. Furthermore, a ZT framework for e-health environments is proposed in [98] based on user trust levels and sensitivity of resources. The authors defined a trust score that ranges from 0 to 100, where 0 is no trust and 100 represents complete trust, and sensitivity score that bears the same value range. Access to a resource is granted, or denied, or may require re-authentication based on the trust level of the user and the sensitivity of the resource requested. In this system, user trust may be reduced by the system if it identifies any access abnormalities through three analysis mechanisms - context-based trust which evaluates how the user interacts with the system, device trust which assesses the device used for access, and historical trust which analyzes past user behavior. The effectiveness of the model was demonstrated through four simulated scenarios.
@@
With the increasing drive toward digital transformation, patient health records have been progressively digitized which enables easy administration and exchange of patient’s records where it can be stored, retrieved, or modified seamlessly [101]. However, this has exposed the records to various security challenges which require adequate security solutions [102]. On this note, Sultana et al. [94] utilized the combined advantages of blockchain immutability and the added security of ZTA and proposed a framework for secure transfer and storage of medical data and image files in an electronic healthcare system. In their work, they highlighted that blockchain ensure the integrity of data while ZTA provides encryption and ensure that only authenticated devices and users access the enterprise network. The proposed system is decentralized to eliminate single point bottleneck, used two-factor authentication to secure access to the proposed system’s web page, smart contract to assign roles to users based on their functions and privileges, while verifying the security status and parameters of any device requesting access to the system.
@@
Furthermore, a ZTA framework integrated into microservice architecture to address security challenges in EHR is proposed in [101]. It leverages several ZTA principles which highlights the importance of continuous verification, granular access controls, and a data-centric approach to protect sensitive health information. The integration of ZTA with microservice architecture allows each microservice to enforce its own security policies, creating isolated trust zones within the system, and thus limits the impact of a security breach in one service on other parts of the system. The effectiveness of the framework is demonstrated through a case study which shows significant improvements in security posture while maintaining acceptable system performance and user experience.